OpenStudy (aaronandyson):

Given: 4 sin theta = 3 cos theta Find 1.sin theta 2.cos theta 3.cot^2 theta - cosec^2 theta 4. 4 cos^2 - 3 sin^2 + 2 @ParthKohli

OpenStudy (aaronandyson):

@robtobey

OpenStudy (aaronandyson):

@robtobey @ParthKohli @eta

Parth (parthkohli):

Obtain the value of tan(theta) by dividing both sides by cos(theta). Then construct a triangle and find the values of sin(theta), cos(theta), cot(theta) and cosec(theta).

OpenStudy (aaronandyson):

So what 4 tan (theta) = 2(theta) @ParthKohli

Parth (parthkohli):

Huh...

OpenStudy (aaronandyson):

?

OpenStudy (aaronandyson):

sorry 4 tan (theta) = 3(theta) @ParthKohli

Parth (parthkohli):

No... show your work.

OpenStudy (aaronandyson):

sin/cos = tan cos/cos = 1 so 4 sin/cos = 2 cos/cos

OpenStudy (aaronandyson):

ugh 4 sin/cos = 3 cos/cos

Parth (parthkohli):

2 sin theta = 3 cos theta how did you get the 4?

OpenStudy (aaronandyson):

My keyboard is lagging its suppose to be 4 sin theta and not 2 sorry for the error.

OpenStudy (aaronandyson):

So what next @ParthKohli

OpenStudy (aaronandyson):

@ParthKohli

OpenStudy (aaronandyson):

@ParthKohli @robtobey

OpenStudy (aaronandyson):

@ribhu

Parth (parthkohli):

Oh, OK. That makes things simpler.

Parth (parthkohli):

Use the identity \(\sin\theta/\cos \theta = \tan \theta.\)

OpenStudy (aaronandyson):

I did and I got tan = 3/4

OpenStudy (aaronandyson):

@ParthKohli

Parth (parthkohli):

Now make a triangle with opposite = 3 units and adjacent = 4 units.

OpenStudy (aaronandyson):

Sorry but it's 4 sin theta and not two

OpenStudy (aaronandyson):



OpenStudy (aaronandyson):

@ParthKohli

OpenStudy (aaronandyson):

@ParthKohli

OpenStudy (aaronandyson):

@ribhu

OpenStudy (aaronandyson):

@ParthKohli

OpenStudy (aaronandyson):

sin = 3/5 cos = 4/5 right?

Parth (parthkohli):

well done

OpenStudy (aaronandyson):

Now?

OpenStudy (aaronandyson):

3.cot^2 theta - cosec^2 theta

OpenStudy (aaronandyson):

@ParthKohli

Parth (parthkohli):

Just plug in all values for the expressions.

Parth (parthkohli):

\(\cot\theta = 1/\tan \theta\), \(\csc\theta = 1/\sin\theta.\)

OpenStudy (aaronandyson):

tan is 3/4??

OpenStudy (aaronandyson):

or 4/3???

OpenStudy (aaronandyson):

@ParthKohli

Parth (parthkohli):

3/4.

OpenStudy (aaronandyson):

so cot is 4/3??

Parth (parthkohli):

Correct.

OpenStudy (aaronandyson):

cos = 5/3??

Parth (parthkohli):

How can cosine exceed 1?

OpenStudy (aaronandyson):

?

OpenStudy (aaronandyson):

What?

OpenStudy (aaronandyson):

I used the expression you have me.

Parth (parthkohli):

Then you mean cosec theta = 5/3, not cos theta.

OpenStudy (aaronandyson):

Sorry my keyboard is lagging . Help me more @ParthKohli ??

OpenStudy (aaronandyson):

4. 4 cos^2 - 3 sin^2 + 2?? @ParthKohli

Parth (parthkohli):

Plug in the values of cos theta and sin theta into the expression.

OpenStudy (aaronandyson):

@ParthKohli

Parth (parthkohli):

You've tagged me 16 times...

Parth (parthkohli):

Plug in the values of sin theta and cos theta into the expression

OpenStudy (aaronandyson):

35/25??

OpenStudy (aaronandyson):

@poopsiedoodle .....